NURBS 2-4-99 G. Farin

Programming assignment #1

due 2-13

Using the postscript routines from the ftp

site pub/farin/ratbook/ps on bachus.eas.asu.edu, write

a program that draws several copies of your initials in the circular model. Some of the copies should cross the horizon.

Use the program maincirc.ps. A sample input file is circ.dat. It will generate an output file circ.ps. Then modify the input file accordingly.

Here it comes, with some comments:

2: how many lines (”howmany” in program)

1 0 50 0: just use these – they are plotting parameters

0 0 1: projective coordinates of first point 10 10 1: of second

1 0 50 0: same as above for second line

10 10 1 0 10 1